# Qingyu Shi1  Lu Qi2  Jianzong Wu1   Jinbin Bai3

### v2

@@ 6.1 More Implementation Details
Local Image Encoder’s Implementation Details. To enhance region-language alignment of dense features, we employ self-distillation on CLIP-ViT-bigG. The training is conducted on the train2017 split of the COCO dataset for six epochs, using 8 A100 GPUs with a batch size of 2 per GPU. We apply the Adam optimizer with a learning rate of $51e-$ 5 and a weight decay of 0.1. The Local Image Encoder, containing 1.8B parameters, extracts local tokens that are concatenated with image tokens during fine-tuning and inference to improve the alignment of object appearance. This distillation process enhances the performance of our base model, MS-Diffusion, in multi-object generation tasks. However, significant object confusion still occurs during relation generation.
@@
Baselines’ Implementation Details. For tuning-based methods, we implement Textual Inversion, DreamBooth, and Custom Diffusion using their respective Diffusers versions, with learning rates and tuning steps aligned to those reported in the original papers. We utilize the official implementations and checkpoints for training-based methods, adjusting hyperparameters as needed during evaluation. Specifically, we set the scale to 0.6 in MS-Diffusion and sample 30 steps using the EulerDiscreteScheduler. For the SSR Encoder, we employ the UniPCMultistepScheduler, sampling 30 steps and adjusting the scale for each object to accommodate different cases. For $\lambda$-Eclipse, we apply the default settings of the official implementation without modification.
6.2 RelationBench
In this section, we show the object and text prompt contained in our RelationBench in Fig. 10.
@@ 6.2 RelationBench
1 A { } is playing guitar on a park bench, serenading passersby.
2 A { } is playing piano in a grand hall filled with glowing chandeliers.
3 A { } is eating dinner in a bustling restaurant with soft jazz playing in the background.
4 A { } is dancing in the moonlight on a rooftop terrace with a stunning city view.
5 A { } is lifting weights in a modern gym, pushing through an intense workout.
6 A { } is reading a book by the fireplace on a rainy evening.
@@
9 A { } is cooking lunch in an open-air kitchen overlooking a lush green valley.
10 A { } is singing on stage during a vibrant music festival with thousands of fans.
11 A { } is riding a bike along the scenic countryside road during sunset.
12 A { } is riding a horse through a dense forest, surrounded by nature.
13 A { } is riding a motorbike on a winding mountain road, feeling the wind rush by.
14 A { } is playing soccer on a sandy beach, with a vibrant sunset in the background.
15 A { } is playing chess with a { } in a quiet park under the shade of a tree.
16 A { } is partner dancing with a { } in a vintage ballroom with live music playing.
17 A { } is carrying a { } across a rushing river, carefully finding their steps.
18 A { } is fencing with a { } in an elegant arena, with spectators watching closely.
19 A { } shakes hands with a { } at a formal awards ceremony after receiving a prize.
20 A { } is kissing a { } in the rain under a large umbrella in a romantic city square.
@@
6.3 More Results of Relation-aware Image Customization
More Results of Single-object relation-aware generation. We show more qualitative results in Fig. 12. We compare our methods with both training-based and tuning-based methods.
@@ 6.3 More Results of Relation-aware Image Customization
More Results of Multi-object relation-aware generation. In this section, we present additional results on multi-object relation-aware generation. In addition to implementing our RelationBooth on MS-Diffusion, we also use SDXL as a base model. As shown in Fig. 15, our RelationBooth enhances SDXL’s ability to generate images that accurately depict object relations. Compared to ReVersion (Huang et al., 2023), our method generates more precise relations, without object confusion or missing objects, demonstrating the strong performance of RelationBooth in the Relation Generation task.
Figure 15: Our RelationBooth is compatible with SDXL to address Relation Inversion task.
Failure cases. As shown in Fig. 16, we present three typical failure cases from our experimental results. The first involves unreasonable relation generation requests, such as asking a plushie octopus, which lacks limbs, to ’shake hands.’ In response, our model generates additional arms for the plushie octopus, leading to a mismatched appearance. The second issue is the unnaturalness of some generated relations, such as a duck failing to make contact with a cat as it should. The final failure case is object confusion at the interaction point, which is a common challenge across all multi-object generation models.
Figure 16: Failure cases of our RelationBooth.
6.4 Additional Analysis on Our Proposed Components
Our well-curated Data. Our curated data are generated by DALLE-3 (Betker et al., 2023), which can effectively preserve the identity of common categories, such as animals, through its multi-turn dialogue capability. While our fine-tuning dataset consists solely of animals, our method focuses primarily on the relation information in the ground truth by disentangling identity using an off-the-shelf identity extractor. Therefore, the specific categories in the fine-tuning data are not the most critical factor for relation learning.
Figure 17: Our fine-tuning dataset as an example.
Keypoint Matching Loss. We use X-Pose as our keypoint detector due to its open-vocabulary detection capabilities. The keypoint matching loss (KML) facilitates relation generation by explicitly guiding the model’s pose manipulation, resulting in more accurate pose generation. As shown in Fig. 19, the cat’s arm crosses the dog’s body, accurately depicting the ’hug’ relation.
Figure 18: Illustration of Keypoint Matching Loss
Figure 19: Additional ablation studies on KML and Local Tokens.
Local Tokens Visualization. To understand why local features enhance relation-aware generation, we adopt Principal Component Analysis (PCA) to project the dense feature more compactly. As shown in Fig. 20, dense features provide more fine-grained information than CLIP image tokens, aiding in distinguishing between different objects during the generation process and helping to avoid object confusion, especially in cases of heavy overlap. Moreover, it can facilitate object appearance alignment.
6.5 Social Compact
@@
6.6 Incorporate with CogVideoX-5b-I2V
Figure 21: Incorporate with CogVideoX-5b-I2V.

### v3

@@ 8 Additional Implementation Details
Local Image Encoder’s Implementation Details. To enhance region-language alignment of dense features, we employ self-distillation on CLIP-ViT-bigG. The training is conducted on the train2017 split of the COCO dataset for 6 epochs, using 8 A100 GPUs with a batch size of 2 per GPU. We apply the Adam optimizer with a learning rate of $51e-$ 5 and a weight decay of $0.1$. The Local Image Encoder, containing 1.8B parameters, extracts local tokens that are concatenated with image tokens during fine-tuning and inference to mitigate the confusion problem between objects.
@@
Baselines’ Implementation Details. We incorporate ReVersion [11] with MS-Diffusion [35] by its official implementation. We fine-tuning a learnable text embedding on a set of relation-specific images. We inject the text embedding into text prompt embedding during inference. We implement ReVersion [11] on DreamBooth [31] by similar procedure. For tuning-based methods, we implement Textual Inversion [7], DreamBooth [31], and Custom Diffusion [15] using their respective diffuser versions, with learning rates and tuning steps aligned to those reported in the original papers. We implement Mix-of-Show [8] by their official repository. We utilize the official implementations and checkpoints for training-based methods, adjusting hyperparameters as needed during evaluation. Specifically, we set the scale to 0.6 in MS-Diffusion [35] and sample 30 steps using the EulerDiscreteScheduler. For the SSR Encoder [43], we employ the UniPCMultistepScheduler, sampling 30 steps and adjusting the scale for each object to accommodate different cases. For $\lambda$-Eclipse, we apply the default settings of the official implementation without modification.
9 Failure cases
As shown in Fig. 18, we present three typical failure cases from our experimental results. The first involves unreasonable relation generation requests, such as asking a plushie octopus, which lacks limbs, to ’shake hands.’ In response, our model generates additional arms for the plushie octopus, leading to a mismatched appearance. The second issue is the unnaturalness of some generated relations, such as a duck failing to make contact with a cat as it should. The final failure case is object confusion at the interaction point, which is a common challenge across all multi-object generation models.
Figure 18: Failure cases of our DreamRelation.
10 RelationBench
In this section, we show the objects and text prompts contained in our RelationBench in Fig. 19, Fig. 20, and Tab. LABEL:tab:catpion. The objects are selected from well-known benchmarks DreamBench and CustomConcept101, covering the commonly seen categories in the real world. The relations in text prompts have covered the most common relations in the real world.
@@ 10 RelationBench
1 A { } is playing guitar on a park bench.
2 A { } is playing piano in a grand hall.
3 A { } is eating dinner in a bustling restaurant.
4 A { } is dancing in the moonlight.
5 A { } is lifting weights in a modern gym.
6 A { } is reading a book by the fireplace.
@@
9 A { } is cooking lunch in a kitchen.
10 A { } is singing on stage during a vibrant music festival.
11 A { } is riding a bike along the scenic countryside road.
12 A { } is riding a horse on the grassland.
13 A { } is riding a motorbike on the street.
14 A { } is playing soccer on football playground.
15 A { } is playing chess with a { } under a tree.
16 A { } is partner dancing with a { } in a vintage ballroom.
17 A { } is carrying a { } on the diving room.
18 A { } is fencing with a { } in an elegant arena.
19 A { } shakes hands with a { } in the forest.
20 A { } is kissing a { }.
@@
11 Incorporate with CogVideoX-5b-I2V
Additionally, we use our generated relation-aware customized images as the first frame to generate videos by the CogVideoX-I2V model, the generated results are shown in Fig. 22. CogVideoX-I2V can handle simple relations such as ”shaking hands” but struggles with complex interactions such as ”hugging”.
Figure 22: Incorporate with CogVideoX-5b-I2V.
12 Social Impact